Write using positive exponents only. b^–2
Found 2 solutions by Nate, stanbon:
Answer by Nate(3500)
Answer by stanbon(57387)

The negative in the exponent means to "invert".
The 2 means to square.
b^-2 = (b^2)^-1 = 1/b^2
or you can reverse the processes and get the same answer:
b^-2 = (b^-1)2 = (1/b)^2 = 1/b^2
===============
Cheers,
Stan H.